Select 3 different digits from these numbers (987621). Use only these numbers as coefficient and create a matrix of 5x5 (with all coefficient non zero) that has rank 3. Also, explain why the rank is equal to 3.




1
Expert's answer
2021-09-29T06:31:49-0400

Let the "5\times 5" matrix made from the digits "1,2,6."

Let "\mathbf{A}=\left(\begin{array}{lllll}
1 & 1 & 1 & 1 & 1 \\
1 & 2 & 2 & 2 & 2 \\
2 & 2 & 2 & 2 & 2 \\
1 & 1 & 6 & 6 & 6 \\
6 & 6 & 6 & 6 & 6
\end{array}\right)"

To calculate matrix rank transform matrix to ụpper triangular form, using elementary row operations.

"\mathrm{R}_{2}-1 \mathrm{R}_{1} \rightarrow \mathrm{R}_{2} (multiply \ 1 \ row \ by \ 1 \ and \ subtract \ it \ from \ 2 \ row) ;\\ \mathrm{R}_{3}-2 \mathrm{R}_{1} \rightarrow \mathrm{R}_{3} (multiply \ 1 \ row \ by \ 2 \ and \ subtract \ it \ from \ 3 \ row);\\ \mathrm{R}_{4}-1 \mathrm{R}_{1} \rightarrow \mathrm{R}_{4} (multiply \ 1 \ row \ by \ 1 \ and \ subtract \ it \ from \ 4 \ row ) ;\\ \mathrm{R}_{5}-6 \mathrm{R}_{1} \rightarrow \mathrm{R}_{5} (multiply \ 1 \ row \ by \ 6 \ and \ subtract \ it \ from \ 5 \ row )"

"\left(\begin{array}{lllll}
1 & 1 & 1 & 1 & 1 \\
0 & 1 & 1 & 1 & 1 \\
0 & 0 & 0 & 0 & 0 \\
0 & 0 & 5 & 5 & 5 \\
0 & 0 & 0 & 0 & 0
\end{array}\right)\\
 
 \mathrm{R}_{3} \leftrightarrow \ \mathrm{R}_{4} (interchange \ the \ 3 \ and \ 4 \ rows)\\
 
\left(\begin{array}{lllll}
1 & 1 & 1 & 1 & 1 \\
0 & 1 & 1 & 1 & 1 \\
0 & 0 & 5 & 5 & 5 \\
0 & 0 & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & 0
\end{array}\right)\\"

"\begin{aligned}
&\mathrm{R}_{3} / 5 \rightarrow \mathrm{R}_{3} \text { (divide the } 3 \text { row by } 5 \text { ) } \\
&\left(\begin{array}{lllll}
1 & 1 & 1 & 1 & 1 \\
0 & 1 & 1 & 1 & 1 \\
0 & 0 & 1 & 1 & 1 \\
0 & 0 & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & 0
\end{array}\right)
\end{aligned}"

Answer. Since there is 3 non-zero rows, then "\operatorname{Rank}(\mathbf{A})=3 ."